# SAMG: State-Action-Aware Offline-to-Online Reinforcement Learning with Offline Model Guidance

### v1

@@ Abstract
The offline-to-online (O2O) paradigm in reinforcement learning (RL) utilizes pre-trained models on offline datasets for subsequent online fine-tuning. However, conventional O2O RL algorithms typically require maintaining and retraining the large offline datasets to mitigate the effects of out-of-distribution (OOD) data, which limits their efficiency in exploiting online samples. To address this challenge, we introduce a new paradigm called SAMG: State-Action-Conditional Offline-to-Online Reinforcement Learning with Offline Model Guidance. In particular, rather than directly training on offline data, SAMG freezes the pre-trained offline critic to provide offline values for each state-action pair to deliver compact offline information. This framework eliminates the need for retraining with offline data by freezing and leveraging these values of the offline model. These are then incorporated with the online target critic using a Bellman equation weighted by a policy state-action-aware coefficient. This coefficient, derived from a conditional variational auto-encoder (C-VAE), aims to capture the reliability of the offline data on a state-action level. SAMG could be easily integrated with existing Q-function based O2O RL algorithms. Theoretical analysis shows good optimality and lower estimation error of SAMG. Empirical evaluations demonstrate that SAMG outperforms four state-of-the-art O2O RL algorithms in the D4RL benchmark.
Introduction
@@
Offline reinforcement learning (RL)(Lowrey et al. 2019; Fujimoto, Meger, and Precup 2019; Mao et al. 2022; Rafailov et al. 2023) has gained significant popularity due to its isolation from online environments. It relies exclusively on the offline datasets, which can be generated by one or several policies, constructed from historical data or even generated randomly. This paradigm eliminates the risks and costs associated with online interactions, thus offering a safe and efficient pathway to pre-train well-behaved RL algorithms. However, existing offline RL algorithms exhibit an inherent limitation that the offline dataset only covers a partial distribution of the state-action space(Prudencio, Maximo, and Colombini 2023). Therefore, standard online RL algorithms fail to resist the cumulative overestimation on state-action pairs out of the offline distribution(Nakamoto et al. 2023). To this end, most offline RL algorithms limit the decision-making scope of the estimated policy within the offline dataset distribution(Kumar et al. 2019; Yu et al. 2021; Janner et al. 2019). Accordingly, offline RL algorithms are confined in performance by the offline dataset distribution.
@@
To overcome the performance limitations of offline RL algorithms and further improve their performance, it is inspiring to perform an online fine-tuning process with the offline pre-trained model. Similar to the successful paradigm of fine-tuning in deep learning(Weiss, Khoshgoftaar, and Wang 2016; Iman, Arabnia, and Rasheed 2023), this paradigm, categorized as offline-to-online (O2O) RL algorithms, is anticipated to enable substantially faster convergence compared to online RL. However, online fine-tuning process inevitably encounters out-of-distribution (OOD) samples which are laid aside in offline pre-training process. This leads to another dilemma that the overestimated pre-trained model may be misguided toward structural damage and performance deterioration when coming across OOD samples(Nair et al. 2020; Kostrikov, Nair, and Levine 2022). As a result, the O2O RL algorithms tend to remain unchanged or even sharply decline in the initial stage of the fine-tuning process. Existing O2O RL algorithms conquer this by maintaining access to the offline dataset and training with the offline data to restore offline information and restrict model deterioration.
@@
Specifically, some fine-tuning algorithms directly inherit the offline dataset as the replay buffer and only get access to online data by incrementally replacing offline data with online ones through iterations (Lyu et al. 2022; Lee et al. 2022; Wang et al. 2024). This paradigm is tedious given that the sample size of the offline datasets tends to exceed the order of millions (Fu et al. 2020). Hence, these algorithms exhibit low inefficiency in leveraging online data. Other algorithms adopt hybrid configurations in which a buffer is evenly composed of both online and offline samples (Song et al. 2023; Nakamoto et al. 2023). Though this setting mitigates the inefficiency, it still visits a considerable amount of offline data and has not departed from the burden of offline data. In summary, existing O2O RL algorithms significantly compromise the efficiency of utilizing online data to mitigate the negative impact of OOD samples.
@@
Figure 1: Compared with the vanilla O2O framework (a), the proposed SAMG (b) uses a frozen offline RL model with an adaptive distribution-aware model to achieve online fine-tuning with 100% online sample rates.
This compromise could result in several undesirable outcomes. Firstly, the quality of offline data is often heterogeneous, with portions of the data potentially being non-conducive to algorithmic improvement. Consequently, the effective information density within the offline dataset is relatively low. Secondly, the inefficiency in accessing online samples leads to a reduced capacity for exploiting new online knowledge, thereby complicating the process of model enhancement. This inefficiency contradicts the primary objective of the fine-tuning process, which aims to refine the offline model by integrating and absorbing novel information. Hence, existing O2O RL algorithms are hindered in both performance and efficiency.
To tackle the dilemma of low online sample utilization, it is constructive to directly leverage the offline critic, which is extracted from the offline dataset and represents an abstraction of the offline information. To this end, a novel online fine-tuning paradigm is proposed in this paper, named SAMG: State-Action-aware offline-to-online reinforcement learning with offline Model Guidance, which eliminates the burden of offline data and achieves 100% online sample utilization. SAMG freezes the offline critic which contains the offline cognition of the values given a state-action pair and offers guidance for online updating. SAMG combines the offline and online critic weighted by a state-action-aware coefficient to give a compound perspective of comprehension. The state-action-aware coefficient represents a class of functions that denotes the offline confidence of a given state-action pair. SAMG could be easily implemented based on Q-learning based RL algorithms, thus demonstrating strong applicability.
@@
The main contributions of this paper are summarized as follows: (1) the compact offline information generated by the offline model is integrated into the fine-tuning process to guide the updating, (2) a novel class of state-action-adaptive functions is designed to provide distribution-aware weight. (3)Rigorous theoretical proofs demonstrate the convergence and estimation error. (4) by integrating SAMG into four Q-learning based algorithms, we evaluate the performance of four vanilla state-of-the-art offline RL and their corresponding SAMG-enhanced algorithms on the D4RL benchmark tasks, demonstrating the significant advantages of our algorithms.
Related works
Online RL with assistance of dataset. Expert level demonstrations are considered and combined with online data to accelerate online training process (Hester et al. 2018; Li et al. 2023; Nair et al. 2018; Rajeswaran et al. 2018; Vecerik et al. 2017), while other offline data is utilized to insert into the online replay buffer to supplement online learning process, without the requirement of being expert level, whose setting is called hybrid RL (Hester et al. 2018; Schaal 1996; Song et al. 2023). Auxiliary behavioral cloning losses with policy gradients are utilized to guide the updates and accelerate convergence (Kang, Jie, and Feng 2018; Zhu et al. 2019, 2018).
O2O RL. Large task-agnostic datasets are utilized to pre-train a model offline and feed for subsequent RL tasks online (Aytar et al. 2018; Baker et al. 2022; Fujimoto, Meger, and Precup 2019; Lifshitz et al. 2024; Yuan et al. 2024). Recent developments in model-based algorithms address high-dimensional problems and facilitate smooth, efficient adaptation to online samples (Lowrey et al. 2019; Mao et al. 2022; Rafailov et al. 2023).
A different category of work is to train the offline models with provided datasets and then fine-tune the pre-trained models online based on the offline datasets (Beeson and Montana 2022; Kostrikov, Nair, and Levine 2022; Lee et al. 2022; Lyu et al. 2022; Mark et al. 2022; Nair et al. 2020; Nakamoto et al. 2023; Wang et al. 2024; Wu et al. 2022). This line of work is based on offline algorithms (Kumar et al. 2020; Levine et al. 2020; Nair et al. 2020) and falls into the trap of buffer dependency. Most of these works inherit the large offline buffer while Cal-QL (Nakamoto et al. 2023) adopts the hybrid RL setting by evenly mixing online and offline samples (Song et al. 2023). Nevertheless, we demonstrate that they both reduce sample efficiency and struggle to handle changed environment dilemmas.
The pipeline of our work is most closely aligned with the approach proposed by Wang et al.(Wang et al. 2024), in which each paradigm is applicable to RL algorithms utilizing a Q-function. However, our specific implementation diverges significantly. Our method mainly differs from previous ones (Beeson and Montana 2022; Kostrikov, Nair, and Levine 2022; Lee et al. 2022; Lyu et al. 2022; Mark et al. 2022; Nair et al. 2020; Wu et al. 2022) that does not retain offline buffer and directly utilize the offline model for assistance instead.
Preliminaries
@@
RL problem is defined as a sequential decision-making process, where a RL agent interacts with an environment in the form of Markov Decision Process (MDP), which is formulated as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma,\tau)$. $\mathcal{S}$ represents the state space and $\mathcal{A}$ represents the action space. $P(s^{\prime}|s,a)$ denotes the unknown function of dynamics model and $r(s,a)$ denotes the unknown function of reward model limited by $|r(s,a)|\leq R_{max}$. $\gamma\in(0,1)$ denotes the discount factor for future reward and $\tau$ denotes the initial state distribution. The goal of RL is to acquire a policy $\pi(a|s)$ to maximize the cumulative discounted value function or state-value function, which are defined as $V^{\pi}(s)={\mathbb{E}_{a_{k}\sim\pi(s_{k})}[\sum_{k}\gamma^{k}r(s_{k},a_{k})|s_{0}=s]}$ and ${Q^{\pi}(s,a)={\mathbb{E}_{a_{k}\sim\pi(s_{k})}\left[\sum_{k}\gamma^{k}r(s_{k},a_{k})|s_{0}=s,a_{0}=a\right]}}$ respectively.
For actor-critic based RL algorithms, the training process alternates between policy evaluation and policy improvement. Policy evaluation phase maintains an estimated Q-function $Q_{\theta}(s,a)$ parameterized by $\theta$ and updates by applying the Bellman operator, $\mathcal{B}^{\pi}Q\doteq r+\gamma P^{\pi}Q$ (referred to as standard Bellman operator), where $P^{\pi_{\phi}}$ couples the dynamics function with the policy: $P^{\pi_{\phi}}Q(s,a)=\mathbb{E}_{s^{\prime}\sim P(s^{\prime}|s,a),a^{\prime}\sim\pi_{\phi}(a^{\prime}|s^{\prime})}[Q(s^{\prime},a^{\prime})]$. While policy improvement phase sustains a policy $\pi_{\phi}(a|s)$ parameterized by $\phi$ and updates the policy towards higher estimated values by the $Q_{\theta}(s,a)$.
SAMG: methodology
@@
Our approach, Offline-Model-Guided and Distribution-Aware (SAMG) O2O RL, seeks to leverage the pre-trained offline model, rather than relying on the tedious offline data, to guide the online fine-tuning process. Therefore two critical questions arise:
@@
    How can we accurately extract the information contained within the offline model?
@@
    How can we adaptively incorporate the information into the online training process?
We introduce a model-guided technique alongside a state-action-aware weighting mechanism to resolve these issues.
Offline-model-guided technique
Offline-model-guided technique is designed to solve the problem 1. For an algorithm containing a state-action value function $Q_{(}s,a)$, which is approximated by some network $Q_{\theta}(s,a)$ parameterized by $\theta$. This function represents the quality given a specific state-action pair in the view of the offline dataset. So we can freeze and preserve this well-trained offline Q-network and query the offline opinion of the online state-action pairs. To combine the offline information with the online information, we integrate the frozen offline opinion with online Q-values with some weight. Formally, the following policy evaluation equation can be obtained:
$Q(s,a)=$ $r(s,a)+\gamma\left[(1-\alpha(s,a))Q(s^{\prime},a^{\prime})\right.$ (1)
$+\left.{\color[rgb]{0,0,1}{\alpha(s,a)Q^{off}(s^{\prime},a^{\prime})}}\right].$
where the Q-network and $\alpha\in(0,1)$ denote a function class that gives a state-action-adaptive weight and can be implemented with any reasonable form. Novel parts that differ from the standard Bellman equation is marked in blue.
Distribution-aware weight
The function class $\alpha(s,a)$ provides the distribution-aware weights to solve the problem 2. Intuitively, we tend to assign high weights to samples within the offline distribution because we have prior comprehension of these samples and the judgment of the offline model is reliable. As for samples distant from the offline distribution, which can be treated as OOD samples, we have limited knowledge about them and tend to assign low weights to them. Any structure that satisfies this requirement could be treated as an instantiation of SAMG.
In this paper, the conditional variational auto-encoder (C-VAE)  (Kingma et al. 2014) is mainly utilized to infer the offline confidence of online samples, which is broadly used in offline algorithms to represent the behavior policy (Fujimoto, Meger, and Precup 2019; Kumar et al. 2019; Xu, Zhan, and Zhu 2022), and an embedding network (Badia et al. 2019) is briefly introduced as well. Previous works leverage the C-VAE structure to construct the divergence error, which is distant from our setting. While Xu (Xu, Zhan, and Zhu 2022) adopts a similar setting to our work, their C-VAE structure is state-conditional. We find such structure might suffer from posterior collapse (Lucas et al. 2019; Wang, Blei, and Cunningham 2021) in certain environments, as illustrated in Appendix B.1.1.
Posterior collapse implies that the encoder completely fails and the KL-divergence term vanishes to zero for any input. Thus, the decoder structure actually takes noise as input and reconstructs a sample all by itself. This phenomenon is extremely detrimental in RL setting because the output of the encoder is needed for the distribution-aware weight but it now fails to operate.
To overcome this challenge, the C-VAE structure is updated that the encoder takes the state-action pairs as input to meet the state-action-adaptive requirement and the decoder reconstructs from the latent layer conditioned on state-action pairs (similar to the role of RL environments). We also utilize the KL-annealing technique (Bowman et al. 2015) in some environments, detailed in Appendix B.1.2. Through this approach, our state-action-aware C-VAE structure could better resist the posterior collapse. Formally, SAMG adopts the following evidence lower bound (ELBO) on the offline dataset to get the state-action-aware C-VAE:
$\max\limits_{\psi_{1},\psi_{2}}\mathbb{E}_{z\sim\text{Enc}_{\psi_{1}}}$ $\left[\log\text{Dnc}_{\psi_{2}}(s^{\prime}|z,s,a)\right]$ (2)
$-\beta D_{KL}\left[{\text{Enc}_{\psi_{1}}(z|s,a)}||{\text{Dnc}_{\psi_{2}}(z)}\right]$
where $\text{Enc}_{\psi_{1}}(z|s,a)$ and $\text{Dnc}_{\psi_{2}}(s^{\prime}|z,s,a)$ represent the encoder structure and decoder structure respectively; $\text{Dnc}_{\psi_{2}}(z)$ denotes the prior distribution of the encoder; and $D_{KL}\left[p||q\right]$ denotes the Kullback-Leibler divergence. The former error term denotes the reconstruction loss of the encoder and the latter one denotes the Kullback-Leibler divergence between encoder distribution and the prior distribution of $z$.
After fitting the C-VAE network, a well-converged estimator can be obtained which encodes the state-action pair to a nearly standard normal distribution, i.e., the output mean value is around 0 and the output standard value is approximately 1, as shown in Appendix B.1.3. However, it is impractical to directly utilize the information of the latent layer because the latent layer is sampled from an approximate standard normal distribution and the sampling randomness completely overwhelms the latent layer information.
Therefore, we collect the output mean and standard values of all the offline data and fit the mean and standard values with two different normal distributions, denoted as $\mathcal{N}_{M}(a_{mean},b_{mean}),\mathcal{N}_{S}(a_{std},b_{std})$. Then the usual range of the mean and standard values of the offline data can be obtained and employed to judge the offline confidence of online samples based on these two normal distributions.
Formally, the weight value with the mean probability of the latent layer of online samples $(a,b)$ falling within a distance less than $|a-a_{mean}|$ for mean output and $|b-b_{mean}|$ for standard output. The cumulative distribution function $F_{X}(x)$ is leveraged to get the probability of an online sample aligning with the offline distribution. We find that the estimation of the standard value sometimes has measurable errors. Therefore, we only utilize the mean value in practice. Moreover, in cases where the sample diverges notably from the offline distribution, the information about the sample is unknown so the probability is manually set to be 0, with a hyperparameter $p_{om}$ introduced. The eventual equation to calculate the probability $p_{o}$ of a given sample $(a,b)$ is illustrated below:

### v2

# SAMG: Offline-to-Online Reinforcement Learning via State-Action-Conditional Offline Model Guidance

@@ Abstract
Offline-to-online (O2O) reinforcement learning (RL) pre-trains models on offline data and refines policies through online fine-tuning. However, existing O2O RL algorithms typically require maintaining the tedious offline datasets to mitigate the effects of out-of-distribution (OOD) data, which significantly limits their efficiency in exploiting online samples. To address this deficiency, we introduce a new paradigm for O2O RL called State-Action-Conditional Offline Model Guidance (SAMG). It freezes the pre-trained offline critic to provide compact offline understanding for each state-action sample, thus eliminating the need for retraining on offline data. The frozen offline critic is incorporated with the online target critic weighted by a state-action-adaptive coefficient. This coefficient aims to capture the offline degree of samples at the state-action level, and is updated adaptively during training. In practice, SAMG could be easily integrated with Q-function-based algorithms. Theoretical analysis shows good optimality and lower estimation error. Empirically, SAMG outperforms state-of-the-art O2O RL algorithms on the D4RL benchmark.
Machine Learning, ICML
1 Introduction
@@
Offline reinforcement learning (Lowrey et al., 2019; Fujimoto et al., 2019; Mao et al., 2022; Rafailov et al., 2023) has gained significant popularity due to its isolation from online environments. It relies exclusively on offline datasets, which can be generated by one or several policies, constructed from historical data, or even generated randomly. This paradigm eliminates the risks and costs associated with online interactions and offers a safe and efficient pathway to pre-train well-behaved RL agents. However, offline RL algorithms exhibit an inherent limitation in that the offline dataset only covers a partial distribution of the state-action space (Prudencio et al., 2023). Therefore, standard online RL algorithms fail to resist the cumulative overestimation on samples out of the offline distribution (Nakamoto et al., 2023). To this end, most offline RL algorithms limit the decision-making scope of the estimated policy within the offline dataset distribution (Kumar et al., 2019; Yu et al., 2021). Accordingly, offline RL algorithms are conservative and are confined in performance by the limited distribution.
@@
To overcome the performance limitation of offline RL algorithms and further improve their performance, it is inspiring to perform an online fine-tuning process with the offline pre-trained model. Similar to the successful paradigm of transfer learning in deep learning (Weiss et al., 2016; Iman et al., 2023), this paradigm, categorized as offline-to-online (O2O) RL algorithms, is anticipated to enable substantially faster convergence compared to pure online RL. However, the online fine-tuning process inevitably encounters out-of-distribution (OOD) samples laid aside in the offline pre-training process. This leads to another dilemma: the conservative pre-trained model may be misguided toward structural damage and performance deterioration when coming across OOD samples (Nair et al., 2020; Kostrikov et al., 2022). Therefore, O2O RL algorithms tend to remain unchanged or even sharply decline in the initial stage of the fine-tuning process. Existing O2O RL algorithms conquer this by maintaining access to the offline dataset and retraining on the offline data during online iterations to restore offline information and restrict OOD deterioration.
@@
Specifically, most fine-tuning algorithms directly inherit the offline dataset as online replay buffer and only get access to online data by incrementally replacing offline data with online ones through iterations (Lyu et al., 2022; Lee et al., 2022; Wen et al., 2024; Wang et al., 2024). This paradigm is tedious given that the sample size of the offline datasets tends to exceed the order of millions (Fu et al., 2020). Hence, these algorithms exhibit low inefficiency in leveraging online data. Other algorithms (Nakamoto et al., 2023; Zheng et al., 2023; Guo et al., 2023; Liu et al., 2024) maintain an online buffer and an offline one and sample from the two replay buffers with hybrid setting (Song et al., 2023) or priority sampling technique. Though these settings mitigate the inefficiency, they still visit a considerable amount of offline data and have not departed from the burden of offline data. In summary, existing O2O RL algorithms severely compromise the efficiency of utilizing online data to mitigate the negative impact of OOD samples.
@@
Figure 1: Compared with the vanilla O2O framework (a), the proposed SAMG (b) uses the frozen offline RL model with an adaptive distribution-aware model to achieve online fine-tuning with 100% online sample rates.
This compromise results in several undesirable outcomes. Training with offline data can potentially hindering algorithmic improvement given the sub-optimal nature of some offline data. Meanwhile, the inefficiency in accessing online samples limits the ability to explore and exploit novel information, making model improvement more challenging. In summary, this setting contradicts the goal of the fine-tuning process to improve algorithm performance with limited training budget.
To tackle the challenge of low online sample utilization, it is inspiring to directly leverage the offline critic, which is meticulously extracted from the offline dataset as an abstraction of the offline information. To this end, this paper introduces a novel online fine-tuning paradigm named State-Action-Conditional Offline Model Guidance (SAMG), which eliminates the reliance on offline data and achieves 100% online sample utilization.
SAMG freezes the offline pre-trained critic, which contains the offline cognition of the values given a state-action pair and offers offline guidance for online fine-tuning process. SAMG combines the offline critic with online target critic weighted by a state-action-conditional coefficient to provide a compound comprehension perspective. The state-action-conditional coefficient represents a class of functions that quantify the offline confidence of a given state-action pair. It is adaptively updated during training to provide accurate probability estimation. SAMG avoids introducing inappropriate intrinsic rewards by leveraging this probability-based mechanism. It does not affect offline algorithms and can be easily deployed on Q-function-based RL algorithms, demonstrating strong applicability.
@@
The main contributions of this paper are summarized as follows: (1) The tedious offline data is eliminated to facilitate more effective online sample utilization. (2) The compact offline information generated by offline model is integrated to provide offline guidance. A novel class of state-action-conditional function is designed and updated to estimate the offline confidence. (3) Rigorous theoretical analysis demonstrates good convergence and lower estimation error. SAMG is integrated into four Q-learning-based algorithms, showcasing remarkable advantages.
2 Preliminaries
@@
Reinforcement Learning task is defined as a sequential decision-making process, where an RL agent interacts with an environment modeled as a Markov Decision Process (MDP): $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma,\tau)$. $\mathcal{S}$ represents the state space and $\mathcal{A}$ represents the action space. $P(s^{\prime}|s,a)$ denotes the unknown function of transition model and $r(s,a)$ denotes the reward model bounded by $|r(s,a)|\leq R_{max}$. $\gamma\in(0,1)$ denotes the discount factor for future reward and $\tau$ denotes the initial state distribution. The goal of the RL agent is to acquire a policy $\pi(a|s)$ to maximize the cumulative discounted reward, defined as value function $V^{\pi}(s)={\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}r(s_{k},a_{k})|s_{0}=s\right]}$ or state-action value function $Q^{\pi}(s,a)={\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}r(s_{k},a_{k})|s_{0}=s,a_{0}=a\right]}$. The training process for actor-critic algorithms alternates between policy evaluation and policy improvement phases. Policy evaluation phase maintains an estimated Q-function $Q_{\theta}(s,a)$ parameterized by $\theta$ and updates it by applying the Bellman operator: $\mathcal{B}^{\pi}Q\doteq r+\gamma P^{\pi}Q$, where $P^{\pi_{\phi}}Q(s,a)=\mathbb{E}_{s^{\prime}\sim P(s^{\prime}|s,a),a^{\prime}\sim\pi_{\phi}(a^{\prime}|s^{\prime})}{\bigl{[}Q(s^{\prime},a^{\prime})\bigr{]}}$. In the policy improvement phase, the policy $\pi_{\phi}(a)$ is parameterized by $\phi$ and updated to achieve higher expected returns.
Conditional Variational Autoencoder (C-VAE) is a generative model designed to capture complex conditional data distributions by incorporating additional information. The encoder maps the input data to the mean $z_{m}$ and variance $z_{v}$ parameters of a Gaussian distribution. Latent vector $z$ is then sampled from this estimated distribution $\mathcal{N}(z_{m},z_{v})$ and then fed to the decoder to reconstruct the data (Kingma et al., 2014). C-VAE is broadly used in offline RL to approximate the behavior policy (Fujimoto et al., 2019; Kumar et al., 2019; Xu et al., 2022) or to estimate the state-action density (Guo et al., 2023).
3 SAMG: Methodology
@@
In this section, we introduce the SAMG paradigm, which leverages the pre-trained offline model to guide the online fine-tuning process without relying on tedious offline data. This approach raises three key questions::
@@
    How can we accurately extract the concentrated information contained within the offline model?
@@
    How can we assess the reliability of this information?
  • 3
    How can we adaptively adjust the level of reliability throughout the training process?
To resolve these challenges, We propose a novel model-guidance technique and introduce an adaptive state-action-conditional coefficient, as illustrated in Fig. 2.
Figure 2: Architecture of SAMG. This figure illustrates the structure of SAMG, highlighting the transition from offline pre-training to online fine-tuning. It outlines key components, including offline critic, VAE model, offline-guidance technique, and adaptive coefficient.
3.1 Offline-model-guidance Paradigm
Offline-model-guidance paradigm is designed to address Problem 1. Intuitively, the offline pre-trained value function $Q_{\theta}^{off}(s,a)$ of an algorithm estimates the quality of a specific state-action pair in the perspective of the offline dataset. This well-trained offline Q-network can be frozen and preserved to provide offline opinion when encountering online state-action pairs. To leverage both offline and online sights, the frozen offline Q-values are integrated with online Q-values weighted by a state-action-conditional coefficient. This approach brings several advantages: it can adaptively utilize the offline information based on its reliability and eliminate introducing undesirable intrinsic rewards, which will be discussed later. Formally, the policy evaluation equation can be obtained as follows: